The profit from a firm’s operation can be $1 million, $4 million, $6 million, or $10 million with probabilities 0.1, 0.3, 0.5, and 0.2, respectively. What is the variance for the firm’s profit?
Answer #1
x P(x) xP(x) x2P(x)
1 0.1000 0.1000 0.1000
4 0.3000 1.2000 4.8000
6 0.5000 3.0000 18.0000
10 0.2000 2.0000 20.0000
total 6.3000 42.9000
E(x) =μ= ΣxP(x) = 6.3000
E(x2) = Σx2P(x) = 42.9000
Var(x)=σ2 = E(x2)-(E(x))2= 3.2100

from above: variance for the firm’s profit =$ 3.21 million